OpenStudy (anonymous):

Let u = <-6, 1>, v = <-5, 2>. Find -4u + 2v. <34, -8> <14, 0> <14, 3> <44, -12>

OpenStudy (jdoe0001):

\(\bf \square <a,b>\implies <\square \cdot a,\square \cdot b> \\ \quad \\ <a,b>+<c,d>\implies <a+c,b+d>\) thus do the scalar multiplication and then add the resulting vectors

OpenStudy (anonymous):

Umm how do I do Scalar multiplication?

OpenStudy (jdoe0001):

\(\bf \square <a,b>\implies <\square \cdot a,\square \cdot b>\leftarrow {\color{brown}{ \textit{scalar multiplication}}} \\ \quad \\ <a,b>+<c,d>\implies <a+c,b+d>\)

OpenStudy (anonymous):

My browser isn't showing what the little boxes in front of <a.b> or <?*a,?*b> are...

OpenStudy (jdoe0001):

is just a box, thus, just means any value

OpenStudy (anonymous):

I'm sorry I don't know what goes there. i'm assuming a=-6, b= 1, c=-5 and d=2 but I don't know what to do with the boxes.

OpenStudy (jdoe0001):

\(\large u = <-6, 1>\qquad v = <-5, 2> \begin{array}{llll} {\color{brown}{ -4}}u\implies <{\color{brown}{ \square }}\cdot -6,{\color{brown}{ \square }}\cdot1> \\ \quad \\ {\color{olive}{ 2}}v\implies <{\color{olive}{ \square }}\cdot -5,{\color{olive}{ \square }}\cdot2> \end{array}\)

OpenStudy (jdoe0001):

what would you get for that?

OpenStudy (anonymous):

Is it <-4*-6*u*1> and <2*-5*v*2> ?

OpenStudy (jdoe0001):

yeap

OpenStudy (jdoe0001):

hmmm wait.. a sec

OpenStudy (jdoe0001):

how did the "u" and "v" get there?

OpenStudy (anonymous):

Ummmm -4u<-- and 2v<---?

OpenStudy (jdoe0001):

so you just distribute the "scalar number" inside the vector, that's a scalar multiplication

OpenStudy (anonymous):

is it 24,-20 ????

OpenStudy (jdoe0001):

hmmm what did you get for -4u and 2v?

OpenStudy (anonymous):

I don't know anymore :( I thought i had distributed them into the vector but then you were confused as to where I got the variables from!

OpenStudy (jdoe0001):

well... you had Is it <-4*-6*u*1> and <2*-5*v*2> ? ? also a scalar multiplication gives you a vector back anyhow

OpenStudy (anonymous):

<-8, 30, uv, 2>?

OpenStudy (jdoe0001):

http://www.youtube.com/watch?v=pNMrYACjHXQ <--- this one works better I think